What is indefinite Integration of (1+tanx)^(squareroot)

Grade:12th pass

2 Answers

Nishant Vora IIT Patna
7 years ago
Take 1 + tanx = t
hence sec^2 x dx = dt
sec^2x = 1 + (t-1)^2

so dx = dt /1 + (t-1)^2

So

Integration of (1+tanx)^(squareroot) = root(t) dt /1 + (t-1)^2
Now proceed

Thanks
kamalesh bagrecha
7 years ago
That much i have already done...Tell me after that how ti procede And how could i know that which function is integrable